# Non-linear degenerate parabolic flow equations and a finer differential structure on Wasserstein spaces

### v1

@@ 2.2.3 Functionals on Wasserstein spaces
Let $p\geq 2$ and $n\in\mathbb{N}$ and $\mu\in\mathcal{W}_{p}(M)$. The map
@@
Let $x_{0}\in M$, $\nu\in\mathcal{W}_{p}(M)$ and let $\gamma\in\Gamma^{opt}(\mu,\nu)$ be an optimal transport plan between $\mu$ and $\nu$. Then, we obtain
@@
$\qquad=\biggl{|}\int_{M\times M}(d(x,x_{0})-d(y,x_{0}))(d(x,x_{0})+d(y,x_{0}))\operatorname{d}\gamma(x,y)\biggr{|}$
$\qquad\leq\biggl{(}\int_{M\times M}(d(x,x_{0})+d(y,x_{0}))^{2}\operatorname{d}\gamma(x,y)\biggr{)}^{\frac{1}{2}}\cdot$
$\qquad\qquad\qquad\qquad\qquad\qquad\biggl{(}\int_{M\times M}(d(x,x_{0})-d(y,x_{0}))^{2}\operatorname{d}\gamma(x,y)\biggr{)}^{\frac{1}{2}}$
$\qquad\leq\sqrt{2}\bigl{(}\mathcal{M}_{2}(\mu)+\mathcal{M}_{2}(\nu)\bigr{)}^{\frac{1}{2}}\mathcal{W}_{2}(\mu,\nu).$
Let $B>0$. Since by Theorem 2.1, $\mathcal{M}_{2}:\mathcal{W}_{p}(M)\to\mathbb{R}$ is continuous, there exists $\delta>0$ such that for all $\mu\in B(\mu,\delta)$, $\mathcal{M}_{2}(\mu)\leq B$. In particular, for all $\mu\in B(\mu,\delta)$,
@@
$\int_{F}\varphi(x)\operatorname{d}\mu(x)=\varphi(\int_{F}x\operatorname{d}\mu(x)).$ (2.9)
@@ 2.2.4 Compact subsets of Wasserstein spaces
Fix $x_{0}\in M$. A family $(\mu_{i})\in\mathcal{W}_{p}(M)^{I}$ is called tight in the pp–th moment if for all $\epsilon>0$, there exists a compact set $K$ such that for all $i\in I$, $\int_{M\setminus K}d(x,x_{0})^{p}\operatorname{d}\mu_{i}(x)\leq\epsilon$.
@@
Take $x_{0}\in M$. Let $(\mu_{i})_{i\in I}$ be a family of measures such that there exists $B>0$ such that for all $i\in I$, $\mathcal{M}_{p}(\mu_{i})\leq B$. Then for all $R>0$ and all $i\in I$,
@@
Furthermore, $(\mu_{i})_{i\in I}$ is tight in the qq–th moment for all $q<p$. More precisely, for all $R\geq R_{\varepsilon}$, with
@@
$\qquad=\int_{R^{q}}^{\infty}\mu_{i}(d(x_{0},x)^{q}>\lambda)\operatorname{d}\lambda+\int_{0}^{R^{q}}\mu_{i}(M\setminus\overline{B(x_{0},R^{q})})\operatorname{d}\lambda$
$\qquad\leq\int_{R^{q}}^{\infty}\mu_{i}(M\setminus B(x_{0},\lambda^{\frac{1}{q}}))\operatorname{d}\lambda+R^{q}\frac{B^{p}}{R^{p}}$
$\qquad\leq\int_{R^{q}}^{\infty}\frac{B^{p}}{\lambda^{\frac{p}{q}}}\operatorname{d}\lambda+\frac{B^{p}}{R^{p-q}}=\frac{B^{p}}{R^{p-q}}\biggl{(}\frac{1}{\frac{p}{q}-1}+1\biggr{)}.$
@@
Let $x_{0}\in M$. Let $(\mu_{i})\in\mathcal{W}_{p}(M)^{I}$ be tight in the pp–th moment. Then $(\mu_{i})$ is tight. Furthermore, let $(i_{n})_{n\in\mathbb{N}}$ be a sequence in $I$ such that $(\mu_{i_{n}})$ converges in the weak–*–topology. Then, $(\mu_{i_{n}})$ converges with respect to the $\mathcal{W}_{p}$–distance.
@@
$\mu_{i}(K\cup B(x_{0},1)))\leq\varepsilon$ (2.18)
and we conclude by compactness of $K\cup B(x_{0},1))$ that $(\mu_{i})$ is tight.
Now, suppose $(\mu_{i_{n}})$ is a sequence in $(\mu_{i})$ that converges weak-*, hence it converges narrowly to some probability measure $\mu$ because the $(\mu_{i})$ are tight by Prokhorov’s Theorem [7, Theorem 2.1.11]. Let $R>0$ such that $K\subset B(x_{0},R)$. We have $\int_{M\setminus B(x_{0},R)}d(x_{0},x)^{p}-R^{p}\operatorname{d}\mu_{i_{n}}(x)<\varepsilon$ for all $n\in\mathbb{N}$. On the other hand, $d_{R}:x\mapsto\min(R^{p},d(x_{0},x)^{p})$ is bounded and continuous. Hence,
$\lim_{n\to\infty}\int_{M}d_{R}(x)\operatorname{d}\mu_{i_{n}}(x)=\lim_{n\to\infty}\int_{M}d_{R}(x)\operatorname{d}\mu(x).$ (2.19)
@@
But for all $\lambda>0$, $\lim_{n\to\infty}\mu_{i_{n}}(\{d(x_{0},x)^{p}>\lambda\})=\mu(\{d(x_{0},x)^{p}>\lambda\})$ by narrow convergence. Hence, Fatou’s Lemma yields $\mathcal{M}_{p}(\mu)\leq\lim_{n\to\infty}\mathcal{M}_{p}(\mu_{i_{n}})<\infty$ and, thus, a posteriori
@@ 2.3 Riemannian Geometry
Let $(M,g)$ be a finite dimensional Riemannian manifold. Then there exists $n\in\mathbb{N}$ and an isometric embedding $f:(M,g)\to(\mathbb{R}^{n},g_{0})$.
@@ 2.3.1 Extension of smooth vector–fields
Let $(M,g)$ be a Riemannian manifold and $R\subseteq M$ a submanifold. The concepts we are about the use are introduced in Michor [14]. Denote the normal bundle of $R$ in $M$ by $N^{g}R$ and the restriction of the Levi-Cività connection on $M$ to $N^{g}R$ by $\nabla^{g}$. Furthermore, let $s_{0}:R\to N^{g}R$ be the 0– section. Note, that any (k,l)(k,l)–tensor $\Phi$ on $R$ can be canonically extended to a (k,l)(k,l)–tensor $N^{g}\Phi$ on $N^{g}R$ by pointwise push-forward and–or pull-back of $\Phi$ under the horizontal lifts induced by $\nabla$ (and–or their inverses respectively). In particular, if $\varphi\in\mathcal{C}^{\infty}$, then $N^{g}\varphi$ is just the function obtained by constantly extending $\varphi$ on each fiber.
@@
Now take $t\mapsto v_{t}$ such that $t\mapsto|v_{t}(p)|_{g}$ $L^{1}$– integrable on an open interval $I$ of $\mathbb{R}$ for all $p\in R$. Take $x\in M$. We have $\overline{v_{t}}_{x}=\eta(x)\operatorname{d}_{t^{-1}(x)}t(\overline{N^{g}}(v_{t})_{t^{-1}(x)})$. Hence,
@@
However, one can show, that if $R$ is of bounded geometry, i.e., the Riemann curvature is bounded in all derivatives, then Lipschitz tensors fields can be extended to Lipschitz tensors, because of the existence of a partition of unity with uniformly bounded derivatives.
@@ 2.3.2 Smooth approximations of Lipschitz functions
$\|f_{\varepsilon}\|_{W^{2,\infty}}<\frac{K}{\varepsilon}.$
@@
Let $n\geq 1$, $M$ be an nn–dimensional manifold, with or without border, with bounded geometry—or in particular compact—, and $f\in W^{1,\infty}_{loc}(M)$ be 11–Lipschitz. Then there exists a constant $K_{M}>0$ depending only on $n$ and $M$ such that for all $\varepsilon>0$, there exists $f_{\varepsilon}\in W^{2,\infty}(\mathbb{R}^{n})$ 11–Lipschitz with
@@
$\|f_{\varepsilon}\|_{W^{2,\infty}}<\frac{K_{M}}{\varepsilon}.$
@@
The Nash embedding Theorem  yields the existence of $n\geq 1$ and an isometric embedding $i:M\to\mathbb{R}^{n}$.
Now, $f\circ i^{-1}:i(M)\to\mathbb{R}$ is 11–Lipschitz. Since $M$ is of bounded geometry, analogously to the last point of Lemma 2.18, there exists $a\geq 1$ — depending only on $M$ — such that $f\circ i^{-1}$ can be extended to an aa–Lipschitz function $\overline{f}$ defined on $\mathbb{R}^{n}$.
@@
Alternatively, for all $\varepsilon>0$, define the $\mathcal{C}^{\infty}$ function
$\mathcal{G}_{\varepsilon}:M$ $\to\mathbb{R}$ (2.27)
$x_{0}$ $\mapsto\int_{M}e^{-\frac{d(x_{0},y)^{2}}{2\varepsilon^{2}}}\operatorname{d}y.$
Note that, since $M$ is of bounded geometry, for all $R>0$, there exits $U_{R},I_{R}>0$ such that for all $x\in M$, $I_{R}\leq\lambda^{g}(B(x,R))\leq U_{R}$. Hence, there exist $M_{\varepsilon},m_{\varepsilon}>0$ such that $\mathcal{G}_{\varepsilon}(x)\geq m_{\varepsilon}$ for all $x\in M$ and $\|\mathcal{G}\|_{W^{2,\infty}(M)}\leq M_{\varepsilon}$. Therefore, we can prove the result by a convolution argument similar to the proof of Proposition 2.20. $\square$
3 Parabolic structure on the Wasserstein space
3.1 Tangent bundle and vector field probabilities
Let us start with a formal setup and introduce some notations. Let $p>2$, $(M,g)$ be a complete finite dimensional Riemannian manifold and $\mathcal{X}^{2,\infty}(M)$ the space of vector fields on $M$ with $W^{2,\infty}$–regularity, endowed with the $W^{2,\infty}$–norm. Denote for all $x\in M$ the evaluation in $x$ by $\pi_{x}:\mathcal{X}^{2,\infty}(M)\to T_{p}M$, $X\mapsto X_{x}$. Furthermore, for all $X\in\mathcal{X}^{2,\infty}(M)$ and all $t\in[0,T]$, the flow of $X$ at time $t$, denoted by $\rho_{t}^{X}:M\mapsto M$, exists by the Theorem of Picard–Lindelöf.
@@ 3.1 Tangent bundle and vector field probabilities
${\operatorname{T}}^{d}_{\mu}\mathcal{W}_{p}(M)\coloneqq\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M),\|\leavevmode\nobreak\ \|_{W^{2,\infty}(M)}).$ (3.1)
@@
While $\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M))$ is not a vector–space, it is a subset of the space of bounded Borel–measures $\mathcal{M}_{p}(\mathcal{X}^{2,\infty}(M))$.
Definition 3.4 (Vector field probability).
Let $V$ be a probability vector–field, i.e. a map
@@
We call $V$ a vector–field probability.
@@
We only require ${\operatorname{T}}^{d}_{\mu}\mathcal{M}_{p}(M)\subset\mathcal{W}_{p}(\mathcal{X}^{0,\infty}(M)$ for the following definition, where $\mathcal{X}^{0,\infty}(M)$ is the space of vector–fields on $M$ with $W^{0,\infty}$–regularity, endowed with the $W^{0,\infty}$–norm. However, the restriction to $L^{\infty}$–regular vector–fields is necessary to ensure that we work on a metric space. Note that many of the following results have similar proofs if we define $V$ more generally as a map taking values in $\mathcal{M}_{p}(\mathcal{X}^{2,\infty}(M))$. For the sake of simplicity, we will not do so here.
@@
Definition 3.6 (Barycenter of a vector field probability).
Denote for all $p\in M$ and all $\mu\in\mathcal{W}_{p}(M)$,
$\overline{V}[\mu](p)\coloneqq\mathbb{E}((\pi_{p})_{*}V[\mu]).$
Then $p\mapsto\overline{V}[\mu](p)\in\mathcal{X}(M)$ is well-defined and shall be denoted by $\overline{V}[\mu]$. Furthermore, we define $\overline{V}:\mu\mapsto\overline{V}[\mu]$, the barycenter of $V$.
A priori, $\overline{V}$ has very little regularity for more general vector field probabilities with image in $V\in\mathcal{W}_{p}(\mathcal{X}^{0,\infty}(M))$. However, the assumption that $V$ takes values in $\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M))$ ensures that $\overline{V}$ is sufficiently well–behaved. Indeed, we obtain the following result.
@@
Let $V$ be a vector–field probability.
@@
      $\mathcal{W}_{p}((\rho^{\overline{V}[\mu]}_{t})_{*}\mu,(\rho^{\overline{V}[\mu]}_{s})_{*}\mu)\leq|t-s|\mathcal{M}_{1}(V[\mu]),$ (3.3)
    • the directional derivatives of $V[\mu]$ are given by
      $\frac{\operatorname{d}V[\mu]}{\operatorname{d}x_{i}}=\int_{\mathcal{X}(M)}\pi_{x}(\frac{\operatorname{d}X}{\operatorname{d}x_{i}})\operatorname{d}V[\mu](X)$ (3.4)
@@
$\int_{T_{x}M}(x,v)\operatorname{d}(\pi_{x})_{*}V[\mu](x,v)$ $=\int_{\mathcal{X}(M)}\pi_{x}(X)\operatorname{d}V[\mu](X)$ (3.6)
$\leq\int_{\mathcal{X}(M)}\|X\|_{\infty}\operatorname{d}V[\mu](X)<\mathcal{M}_{1}(V[\mu]).$
@@
$\leq\int_{\mathcal{X}(M)}|P^{y}_{x}\pi_{x}(X)-\pi_{y}(X)|_{g_{y}}\operatorname{d}V[\mu](X)$
$\leq\int_{\mathcal{X}(M)}\|X\|_{W^{1,\infty}(M)}d(x,y)\operatorname{d}V[\mu](X)\leq d(x,y)\mathcal{M}_{1}(V[\mu]).$
@@
$\mathcal{W}_{p}(\mu,e^{V}_{t}(\mu))^{p}\leq\int_{M}d(x,\rho^{\overline{V}[\mu]}_{t}(x))^{p}\operatorname{d}\mu\leq t^{p}\mathcal{M}_{1}(V[\mu])^{p}.$
@@
$\mathcal{W}_{p}(\mu,e^{V}_{t}(\mu))\leq t\mathcal{M}_{1}(V[\mu]),$ (3.8)

### v2

@@ 2.2.3 Functionals on Wasserstein spaces
Let $p\geq 2$, $n\in\mathbb{N}$ and $(M,g)$ be a Riemannian manifold. The second moment
$\mathcal{M}_{2}:\mathcal{W}_{p}(M)\to\mathbb{R}_{+}$ (2.9)
is locally Lipschitz with respect to the pp–Wasserstein distance. Furthermore, the map
@@
Let $x_{0}\in M$, $\mu,\nu\in\mathcal{W}_{p}(M)$ and let $\gamma\in\Gamma^{\operatorname{opt}}(\mu,\nu)$ be an optimal transport plan between $\mu$ and $\nu$. Then, we obtain
@@
$\hskip 18.49988pt=\biggl|\int_{M\times M}(d(x,x_{0})-d(y,x_{0}))(d(x,x_{0})+d(y,x_{0}))\operatorname{d}\gamma(x,y)\biggr|$
$\hskip 18.49988pt\leq\biggl(\int_{M\times M}(d(x,x_{0})+d(y,x_{0}))^{2}\operatorname{d}\gamma(x,y)\biggr)^{\frac{1}{2}}$
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\cdot\biggl(\int_{M\times M}(d(x,x_{0})-d(y,x_{0}))^{2}\operatorname{d}\gamma(x,y)\biggr)^{\frac{1}{2}}$
$\hskip 18.49988pt\leq\sqrt{2}\bigl(\mathcal{M}_{2}(\mu)+\mathcal{M}_{2}(\nu)\bigr)^{\frac{1}{2}}\mathcal{W}_{2}(\mu,\nu).$
Let $\rho>0$. Since by Theorem 2.7, $\mathcal{M}_{2}:\mathcal{W}_{p}(M)\to\mathbb{R}$ is continuous, there exists $B>0$ such that for all $\mu\in B(\delta_{x_{0}},\rho)$, $\mathcal{M}_{2}(\mu)\leq B$. In particular, for all $\mu,\nu\in B(\delta_{x_{0}},\rho)$,
@@
$\int_{F}\varphi(x)\operatorname{d}\mu(x)=\varphi\biggl(\int_{F}x\operatorname{d}\mu(x)\biggr).$ (2.13)
@@ 2.2.4 Compact subsets of Wasserstein spaces
Let $(M,g)$ be a metric space. Fix $x_{0}\in M$. A family $(\mu_{i})\in\mathcal{W}_{p}(M)^{I}$ is called tight in the pp–th moment if for all $\epsilon>0$, there exists a compact set $K$ such that for all $i\in I$, $\int_{M\setminus K}d(x,x_{0})^{p}\operatorname{d}\mu_{i}(x)\leq\epsilon$.
@@
Let $(M,d)$ be a metric space. Take $x_{0}\in M$. Let $(\mu_{i})_{i\in I}$ be a family of measures such that there exists $B>0$ such that for all $i\in I$, $\mathcal{M}_{p}(\mu_{i})\leq B^{p}$. Then for all $R>0$ and all $i\in I$,
@@
Furthermore, if $(M,d)$ is a Riemannian manifold, $(\mu_{i})_{i\in I}$ is tight in the qq–th moment for all $q<p$. More precisely, for all $R\geq R_{\varepsilon}$, with
@@
$\hskip 18.49988pt=\int_{R^{q}}^{\infty}\mu_{i}(d(x_{0},x)^{q}>\lambda)\operatorname{d}\lambda+\int_{0}^{R^{q}}\mu_{i}(M\setminus\overline{B(x_{0},R^{q})})\operatorname{d}\lambda$
$\hskip 18.49988pt\leq\int_{R^{q}}^{\infty}\mu_{i}(M\setminus B(x_{0},\lambda^{\frac{1}{q}}))\operatorname{d}\lambda+R^{q}\frac{B^{p}}{R^{p}}$
$\hskip 18.49988pt\leq\int_{R^{q}}^{\infty}\frac{B^{p}}{\lambda^{\frac{p}{q}}}\operatorname{d}\lambda+\frac{B^{p}}{R^{p-q}}=\frac{B^{p}}{R^{p-q}}\biggl(\frac{1}{\frac{p}{q}-1}+1\biggr).$
@@
The following results only hold in proper spaces, like Riemannian manifolds.
@@
Let $(M,g)$ be a Riemannian manifold. Let $x_{0}\in M$. Let $(\mu_{i})\in\mathcal{W}_{p}(M)^{I}$ be tight in the pp–th moment. Then $(\mu_{i})$ is tight. Furthermore, let $(i_{n})_{n\in\mathbb{N}}$ be a sequence in $I$ such that $(\mu_{i_{n}})$ converges in the weak–*–topology. Then, $(\mu_{i_{n}})$ converges with respect to the $\mathcal{W}_{p}$–distance.
@@
$\mu_{i}\bigl(M\setminus(K\cup B(x_{0},1))\bigr)\leq\varepsilon$ (2.22)
and we conclude by compactness of $K\cup B(x_{0},1)$ that $(\mu_{i})$ is tight.
Now, suppose $(\mu_{i_{n}})$ is a sequence in $(\mu_{i})$ that converges weak-*, hence it converges narrowly to some probability measure $\mu$ because the $(\mu_{i})$ are tight by Prokhorov’s Theorem [9, Theorem 2.1.11]. Let $R>0$ such that $K\subset B(x_{0},R)$. We have $\int_{M\setminus B(x_{0},R)}d(x_{0},x)^{p}-R^{p}\operatorname{d}\mu_{i_{n}}(x)<\varepsilon$ for all $n\in\mathbb{N}$. On the other hand, $d_{R}:x\mapsto\min(R^{p},d(x_{0},x)^{p})$ is bounded and continuous. Hence,
$\lim_{n\to\infty}\int_{M}d_{R}(x)\operatorname{d}\mu_{i_{n}}(x)=\int_{M}d_{R}(x)\operatorname{d}\mu(x).$ (2.23)
@@
But for all $\lambda>0$, $\lim_{n\to\infty}\mu_{i_{n}}(\{d(x_{0},x)^{p}>\lambda\})=\mu(\{d(x_{0},x)^{p}>\lambda\})$ by narrow convergence. Hence, we get $\mathcal{M}_{p}(\mu)\leq\lim_{n\to\infty}\mathcal{M}_{p}(\mu_{i_{n}})<\infty$ from Fatou’s Lemma and, thus, a posteriori
@@ 2.3 Riemannian Geometry
Let $(M,g)$ be a finite dimensional Riemannian manifold. Then there exists $n\in\mathbb{N}$ and an isometric embedding $f:(M,g)\to(\mathbb{R}^{n},g^{eucl})$.
@@ 2.3.1 Extension of smooth vector–fields
Let $(M,g)$ be a Riemannian manifold and $R\subseteq M$ a submanifold. The concepts we are about to use are introduced in Michor [19]. Denote the normal bundle of $R$ in $M$ by $N^{g}R$ and the restriction of the Levi-Cività connection on $M$ to $N^{g}R$ by $\nabla^{g}$. Furthermore, let $s_{0}:R\to N^{g}R$ be the 0– section. Note, that any (k,l)(k,l)–tensor $\Phi$ on $R$ can be canonically extended to a (k,l)(k,l)–tensor $N^{g}\Phi$ on $N^{g}R$ by pointwise push-forward and–or pull-back of $\Phi$ under the horizontal lifts induced by $\nabla^{g}$ (and–or their inverses respectively). In particular, if $\varphi\in\mathcal{C}^{\infty}$, then $N^{g}\varphi$ is just the function obtained by constantly extending $\varphi$ on each fiber.
@@
Now take $t\mapsto v_{t}$ such that $t\mapsto|v_{t}(p)|_{g}$ is $L^{1}$– integrable on an open interval $I$ of $\mathbb{R}$ for all $p\in R$. Take $x\in M$.
We have $\overline{v_{t}}_{x}=\eta(x)\operatorname{d}_{t^{-1}(x)}t(\overline{N^{g}}(v_{t})_{t^{-1}(x)})$, hence,
@@
However, one can show, that if $R$ is of bounded geometry, i.e., the Riemann curvature is bounded in all derivatives, then Lipschitz tensor fields can be extended to Lipschitz tensors, because of the existence of a partition of unity with uniformly bounded derivatives.
@@ 2.3.2 Smooth approximations of Lipschitz functions
$\|\nabla^{2}f_{\varepsilon}\|_{\infty}<\frac{K}{\varepsilon}.$
@@
Let $n\geq 1$, $M$ be an nn–dimensional manifold, with or without border, with bounded geometry—or in particular compact—, and $f\in W^{1,\infty}_{loc}(M)$ be 11–Lipschitz. Then there exists a constant $K_{M}>0$ depending only on $n$ and $M$ such that for all $\varepsilon>0$, there exists $f_{\varepsilon}\in W^{2,\infty}_{loc}(\mathbb{R}^{n})$ $K_{M}$–Lipschitz with
@@
$\|\nabla^{2}f_{\varepsilon}\|_{\infty}<\frac{K_{M}}{\varepsilon}.$
@@
for compact manifolds, the Nash embedding Theorem yields the existence of $n\geq 1$ and an isometric embedding $i:M\to\mathbb{R}^{n}$.
Now, $f\circ i^{-1}:i(M)\to\mathbb{R}$ is 11–Lipschitz. Since $M$ is compact, analogously to the last point of Lemma 2.25, there exists $a\geq 1$ — depending only on $M$ — such that $f\circ i^{-1}$ can be extended to an aa–Lipschitz function $\overline{f}$ defined on $\mathbb{R}^{n}$.
@@
If we assume just bounded geometry for $M$, this follows from a local application of Proposition 2.27 using a $\mathcal{C}^{2}$–uniformly bounded partition of unity. $\square$
3 Parabolic differential structures on Wasserstein
spaces
3.1 Tangent bundle and probability vector–fields
Let us start with a formal setup and introduce some notation. Let $p>2$, $(M,g)$ be a complete finite dimensional Riemannian manifold and $\mathcal{X}^{2,\infty}(M)$ the space of vector–fields on $M$ with $W^{2,\infty}$–regularity, endowed with the $W^{2,\infty}$–norm. We will always assume completeness of the Riemannian manifold we work with in what follows. Denote for all $x\in M$ the evaluation in $x$ by $\pi_{x}:\mathcal{X}^{2,\infty}(M)\to T_{x}M$, $X\mapsto X_{x}$. Furthermore, for all $X\in\mathcal{X}^{2,\infty}(M)$ and all $t\in[0,T]$, the flow of $X$ at time $t$, denoted by $\rho_{t}^{X}:M\mapsto M$, exists by the Theorem of Picard–Lindelöf.
@@ 3.1 Tangent bundle and probability vector–fields
${\operatorname{T}}^{d}_{\mu}\mathcal{W}_{p}(M)\coloneqq\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M),\|\penalty 10000\ \|_{W^{2,\infty}(M)}).$ (3.1)
The elements of ${\operatorname{T}}^{d}_{\mu}\mathcal{W}_{p}(M)$ will be referred to as tangent vectors at $\mu$ or vector–field probabilites.
Remark 3.2.
The space $\mathcal{X}^{2,\infty}(M)$ is independent of the choice of the atlas on $M$ if $(M,g)$ is of bounded geometry. This is not true for a general Riemannian manifold $M$. However, whenever we do not assume $(M,g)$ to be of bounded geometry, we will always implicitly fix an atlas on $M$. In particular, for those results we prove on more general Riemannian manifolds, one only has to find one atlas such that all tangent vectors considered are in $\mathcal{X}^{2,\infty}(M)$ with respect to this atlas.
@@
While $\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M))$ is not a vector–space, it is a subset of the space of bounded signed Borel–measures $\mathcal{M}_{p}(\mathcal{X}^{2,\infty}(M))$.
Remark 3.5.
Note that $\mathcal{X}^{2,\infty}(M)$ is not a Polish space. Hence, many standard results do not hold on $\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M))$, like the existence of optimal transport plans or the strong Monge–Kantorovich duality of Definition 2.9. However, we shall not need those results for measures on $\mathcal{X}^{2,\infty}(M)$ in this work.
Definition 3.6 (Probability vector–field).
Let $V$ be a map
@@
We call $V$ a probability vector–field.
@@
We only require ${\operatorname{T}}^{d}_{\mu}\mathcal{M}_{p}(M)\subset\mathcal{W}_{p}(\mathcal{X}^{0,\infty}(M))$ for the following definition, where $\mathcal{X}^{0,\infty}(M)$ is the space of vector–fields on $M$ with $W^{0,\infty}$–regularity, endowed with the $W^{0,\infty}$–norm. However, the restriction to $L^{\infty}$–regular vector–fields is necessary to ensure that we work on a metric space. Note that many of the following results have similar proofs if we define $V$ more generally as a map taking values in $\mathcal{M}_{p}(\mathcal{X}^{2,\infty}(M))$. For the sake of simplicity, we will not do so here.
@@
Definition 3.8 (Barycenter of a probability vector–field).
Denote for all $x\in M$ and all $\mu\in\mathcal{W}_{p}(M)$,
$\overline{V}[\mu](x)\coloneqq\mathbb{E}((\pi_{x})_{*}V[\mu]).$
Then $x\mapsto\overline{V}[\mu](x)$ is a well-defined vector field on $M$ and shall be denoted by $\overline{V}[\mu]$. Furthermore, we define $\overline{V}:\mu\mapsto\overline{V}[\mu]$, the barycenter of $V$.
A priori, $\overline{V}$ has very little regularity for more general probability vector–fields with image in $V\in\mathcal{W}_{p}(\mathcal{X}^{0,\infty}(M))$. However, the assumption that $V$ takes values in $\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M))$ ensures that $\overline{V}$ is sufficiently well–behaved. Indeed, we obtain the following result.
@@
Let $V$ be a probability vector–field.
@@
      $\mathcal{W}_{p}\bigl((\rho^{\overline{V}[\mu]}_{t})_{*}\mu,(\rho^{\overline{V}[\mu]}_{s})_{*}\mu\bigr)\leq|t-s|\mathcal{M}_{1}(V[\mu]),$ (3.3)
    • the directional derivatives of $\overline{V}[\mu]$ are given by
      $\frac{\operatorname{d}\overline{V}[\mu](x)}{\operatorname{d}x_{i}}=\int_{\mathcal{X}^{2,\infty}(M)}\pi_{x}\biggl(\frac{\operatorname{d}X}{\operatorname{d}x_{i}}\biggr)\operatorname{d}V[\mu](X)$ (3.4)
@@
$\int_{T_{x}M}(x,v)\operatorname{d}(\pi_{x})_{*}V[\mu](x,v)$ $=\int_{\mathcal{X}^{2,\infty}(M)}\pi_{x}(X)\operatorname{d}V[\mu](X)$ (3.6)
$\leq\int_{\mathcal{X}^{2,\infty}(M)}\|X\|_{\infty}\operatorname{d}V[\mu](X)<\mathcal{M}_{1}(V[\mu]).$
@@
$\leq\int_{\mathcal{X}^{2,\infty}(M)}|P^{y}_{x}\pi_{x}(X)-\pi_{y}(X)|_{g_{y}}\operatorname{d}V[\mu](X)$
$\leq\int_{\mathcal{X}^{2,\infty}(M)}\|X\|_{W^{1,\infty}(M)}d(x,y)\operatorname{d}V[\mu](X)\leq d(x,y)\mathcal{M}_{1}(V[\mu]).$
@@
$\mathcal{W}_{p}\bigl(\mu,(\rho^{\overline{V}[\mu]}_{t})_{*}\mu\bigr)^{p}\leq\int_{M}d(x,\rho^{\overline{V}[\mu]}_{t}(x))^{p}\operatorname{d}\mu(x)\leq t^{p}\mathcal{M}_{1}(V[\mu])^{p}.$
@@
$\mathcal{W}_{p}\bigl(\mu,(\rho^{\overline{V}[\mu]}_{t})_{*}\mu\bigr)\leq t\mathcal{M}_{1}(V[\mu]),$ (3.8)